# Learning to Rank Critical Road Segments via Heterogeneous Graphs with OD Flow Integration

### v2

@@ 5.3 Experimental results
To provide an intuitive comparison of algorithm performance, Figures 4 - 6 illustrate the correspondence between predicted rankings and ground truth importance scores (IS) across the SY-Net110, SY-Net514, and RD-Net3478 datasets. Overall, as predicted rankings decrease, the ground truth IS values for all methods exhibit the expected monotonic downward trend. In the two simulated real-world road networks, HetGL2R demonstrates the smoothest and most stable decline, with minimal scatter, closely matching the ground truth and showing excellent ranking consistency. MGL2Rank, SeHGNN, PDFormer, and Edge-GNN also display reasonable trends, though with slightly larger overall fluctuations. In the randomly generated RD-Net3478 network, HetGL2R maintains robust performance with tightly clustered and orderly scatter points. In contrast, DeepWalk, GAT, GNNLLMEmb, and IPageRank show more dispersed scatter and greater deviation from the true rankings. Structure-driven models such as HGT, GAT, and Dir-GNN perform adequately on the simulated real networks but experience significant performance drops on the randomized network. We attribute this performance difference to the weaker coupling between network topology and OD flows in the synthetic random network. In real-world urban road networks, topology and OD flows are typically strongly coupled: central or highly connected segments are more likely to carry significant OD flows, allowing structural features to serve as reliable proxies for importance. In the RD-Net3478 network, however, this coupling is substantially weaker, making it difficult for topology-dependent models to accurately infer criticality. Furthermore, as a general-purpose heterogeneous graph learning model, HGT is limited by its primary reliance on node and edge types, making it difficult to fully leverage rich edge attributes in path sequences (such as node order), which are crucial for capturing traffic flow directionality and influence scope. HetGL2R serializes the graph structure via heterogeneous random walks and explicitly captures sequential dependencies using the Transformer’s full-sequence modeling capabilities, outperforming message-passing-based heterogeneous GNNs in representing road network structures. Notably, PDFormer achieves high-level performance across all three networks. As a state-of-the-art method designed for traffic networks, it employs an attention mechanism guided by semantic masks to focus on geographically distant but functionally similar node pairs, capturing long-range dependencies, while enhanced key computations simulate propagation delay. Its approach aligns conceptually with HetGL2R in capturing functional and structural dependencies, but differs technically. HetGL2R explicitly models OD flows, paths, and inter-segment relationships to learn more precise long-range dependencies, resulting in more robust performance in experiments. Graph-based large language models have yet to surpass specialized graph learning methods, primarily because they are designed for general graph structures and do not effectively incorporate OD flows, which fundamentally determine long-range spatial dependencies, thereby hindering their ability to accurately capture such dependencies.
@@
In the two simulated real-world road networks, HetGL2R demonstrates the smoothest and most stable decline, with minimal scatter, closely matching the ground truth and showing excellent ranking consistency. MGL2Rank, SeHGNN, PDFormer, and Edge-GNN also display reasonable trends, though with slightly larger overall fluctuations. In the randomly generated RD-Net3478 network, HetGL2R maintains robust performance with tightly clustered and orderly scatter points. In contrast, DeepWalk, GAT, GNNLLMEmb, and IPageRank show more dispersed scatter and greater deviation from the true rankings. Structure-driven models such as HGT, GAT, and Dir-GNN perform adequately on the simulated real networks but experience significant performance drops on the randomized network. We attribute this difference to the weakened correlation between network topology and OD flows in the random network. In real urban road networks, topology and OD flows are typically coupled: central or highly connected segments are more likely to carry major OD flows, allowing structural features to provide strong cues for node importance assessment. In RD-Net3478, however, this topology–OD coupling is weakened, making it difficult for models that rely solely on topology to reliably infer node importance.
Tables 6 to 8 present the predicted rankings of the ground-truth top-10 nodes across the three networks. Comprehensive analysis demonstrates that HetGL2R achieves the highest ranking consistency and PDFormer delivers competitive performance. Both models exhibit exceptional generalization capability and robustness across different network configurations.
Figure 7: Case study on segment ranking in SY-Net514 simulated road network.
Table 9: Runtime and Memory Usage of HetGL2R and Baseline Methods on Three Networks.
@@
Figure 7 illustrates two representative road segments during ranking evaluation on the SY-Net514 network. In the rankings generated by HetGL2R, segment #470 has a relatively low traffic volume (215 vehicles/hour) but is ranked 15th, higher than segment #336, which carries a larger traffic volume (588 vehicles/hour, ranked 22th). Further analysis shows that segment #470 is located near the destination of a certain OD pair, with upstream nodes carrying substantial traffic flow. A failure at segment #470 could block upstream traffic, causing congestion in the upstream region and potentially impacting the broader network performance. This example demonstrates that HetGL2R’s ranking considers not only node attributes or local topology but also the structural dependencies along upstream and downstream paths.
@@ 5.4 Computational Efficiency and Scalability Analysis
To evaluate the scalability of HetGL2R with respect to network size, we analyzed its per-epoch running time on three networks of different scales (Table 9). Our key findings are as follows: (i) Compared to computationally lightweight methods (e.g., DeepWalk, TraNode2vec), specialized ranking models (e.g., LambdaLoss, RankFormer), and the lightweight ensemble model LightGBM, HetGL2R requires significantly longer training time due to its more complex architecture; (ii) However, when compared against other graph neural network models (e.g., PDFormer, DR-HGNN, Edge-GNN), HetGL2R demonstrates comparable efficiency. This comparable efficiency is associated with its random walk sampling strategy. This strategy confines the computational complexity to a linear relationship with the walk length and sampling count, thereby avoiding the rapid increase in computational cost caused by the quadratic or higher-order polynomial complexity with respect to the number of nodes, as seen in methods employing global graph attention (e.g., PDFormer) or deep hierarchical message passing (e.g., DR-HGNN). Regarding scalability, the primary bottlenecks for HetGL2R lie in the Transformer encoder’s processing of long sequences and the storage I/O required for full-graph walk sampling. Consequently, its scalability on very large graphs is still expected to surpass that of most GNNs, yet may not match that of lightweight models requiring no graph structure. In summary, through its core design of trading sampling for computation, HetGL2R achieves an effective balance between the complex requirements of heterogeneous graph ranking and computational feasibility.
@@ 5.5 Ablation Study
Ablation Study on HetGWalk. We evaluated the impact of joint random walks on the Trip Graph (TG) and Attribute-guided Graph (AG) by comparing the full model with a variant excluding AG (w/o AG). The w/o AG variant consistently exhibits substantial performance drops across all networks and metrics, indicating that joint random walks are essential for achieving high ranking accuracy and robustness. This decline arises from TG’s limitations: it captures functional structural dependencies but cannot model functional similarity, and relying solely on topology biases node sampling toward high-centrality nodes while underexploring latent functional clusters. AG mitigates these issues by enabling transitions between semantically similar nodes, enriching contextual diversity. Consequently, the joint TG–AG random walk effectively integrates structural and functional perspectives, generating context-rich sequences that support robust and accurate segment representations.
Ablation on Transformer Encoder. We replaced the Transformer with GRU (gru) and LSTM (lstm) (denoted as w/o transformer (GRU) and w/o transformer (LSTM), respectively). Experimental results show that both RNN variants performed significantly worse than the Transformer architecture across all datasets and metrics. The performance degradation stems from the limitations of RNNs’ sequential modeling mechanism in capturing long-range dependencies, while Transformer’s global self-attention can directly model high-order dependencies between any positions in the sequence, making it more suitable for processing complex sequences generated by heterogeneous graph random walks.
Ablation Study on AMIL aggregation. We performed ablation experiments on the context embedding aggregation strategy by replacing AMIL with Mean, DeepSets (deepsets), and SetTrans (SetTrans) aggregation, denoted as w/o AMIL (Mean), w/o AMIL (DeepSets), and w/o AMIL (SetTrans), respectively. AMIL consistently achieves the best trade-off between performance and efficiency, attaining the highest NDCG@K scores across all datasets and outperforming alternatives on Diff and EMD metrics. The results indicate that Mean fails to differentiate critical contexts, DeepSets lacks inter-instance interaction, and SetTrans overemphasizes fine-grained variations, whereas AMIL’s lightweight attention mechanism effectively captures key contextual information while mitigating noise. These findings validate the effectiveness and superiority of AMIL as the aggregation strategy in our framework.
Ablation study on the KL divergence loss. we replaced the KL divergence with the cross-entropy loss used in ListNet (listnet) and the likelihood loss used in ListMLE (mle), denoted as w/o KL (ListNet) and w/o KL (ListMLE), respectively. Results show that KL divergence outperforms the other losses across all metrics. This reflects the conceptual differences among listwise losses: ListMLE maximizes the likelihood of the true ranking, but its stepwise generation assumption can accumulate errors in noisy traffic network scenarios; ListNet optimizes only the Top-one probability and fails to fully exploit the ranking information of the rest of the list. In contrast, KL divergence performs a soft alignment between the true and predicted importance distributions, with the true distribution (generated by applying Softmax to the true scores) naturally assigning higher weights to the top-ranked nodes, thereby focusing the training on accurate top-ranking predictions. Moreover, KL divergence does not rely on stepwise sequence generation, enabling robust handling of ranking uncertainty and fully utilizing supervision signals from all positions in the list. Based on these properties, KL divergence achieves the best performance in our task.
Figure 8: Parameter sensitivity analysis.
@@
Comprehensive ablation experiments prove that each core component in the HetGL2R framework contributes significantly to the final performance. The organic combination of these components ensures the model’s superior performance in critical road segment identification tasks.
@@ 5.6 Sensitivity analysis
We conducted a sensitivity analysis on key hyperparameters, including the walk bias factor $\alpha$, the sampling sequence length lenlen, the list length $K$, and the depth-first exploration probability $\beta$, with results summarized in Figure 8. When $\alpha$ is set to 0.6, the model achieves the highest Kendall $\tau$, indicating an effective balance between node topology and attribute information. The optimal sampling length lenlen is 20, as shorter sequences fail to capture sufficient contextual information from related nodes, while longer sequences may introduce noise that degrades performance. Similarly, a list length of $K=5$ allows HetGL2R to best capture intra-list ranking dependencies among ranked elements; shorter lists lack internal coherence, whereas longer lists increase task complexity and dilute the ranking signal. Finally, setting $\beta$ to 0.8 yields the highest Kendall $\tau$, showing that a moderately strong preference for depth-first exploration balances exploitation of long-range structural dependencies with control of noise. Values of $\beta$ below 0.6 underutilize multi-hop context and reduce ranking concordance, whereas $\beta=1.0$ effectively switches to pure breadth-first exploration, weakening deep relational signals and sharply decreasing performance. Overall, these results highlight the sensitivity of HetGL2R to these hyperparameters and provide guidance for configuring them to achieve optimal ranking accuracy.
@@ 6 Conclusion and Future Work
This study assumes consistent OD–path–segment mappings derived from accurate OD demand and route choice data. In practice, uncertainties in OD estimation and potential errors in path enumeration may affect the fidelity of the constructed heterogeneous trip graph. Moreover, the current evaluation relies on SUMO-based simulated networks, which provide controlled yet idealized settings. Future work will focus on enhancing robustness to noisy or incomplete OD data, integrating uncertainty modeling, and validating the framework on large-scale real-world traffic networks.

### v3

@@ 5.3 Experimental results
To provide an intuitive comparison of algorithm performance, Figures 4 - 6 illustrate the correspondence between predicted rankings and ground-truth importance scores (IS) across the SY-Net110, SY-Net514, and RD-Net3478 datasets. Overall, as predicted rankings decrease, the ground-truth IS values for all methods exhibit the expected monotonic downward trend. In the two simulated real-world road networks, SY-Net110 and SY-Net514, HetGL2R exhibits the smoothest and most stable decline, with tightly clustered scatter points closely following the ground truth, indicating strong ranking consistency. Several learning-based baselines, including MGL2Rank, SeHGNN, PDFormer, and Edge-GNN, also demonstrate reasonable alignment, albeit with larger fluctuations. In the randomly generated RD-Net3478 network, performance differences among methods become more pronounced. Models that primarily rely on topological structure (e.g., DeepWalk, GAT, GNNLLMEmb, and IPageRank) exhibit more scattered distributions and larger deviations from the true rankings, whereas HetGL2R maintains robust behavior with orderly and concentrated scatter distributions.
@@
Tables 6 to 8 present the predicted rankings of the ground-truth top-10 nodes across the three networks. The results indicate that HetGL2R achieves the highest ranking consistency, while PDFormer delivers competitive performance. Both models demonstrate robust ranking stability across different network configurations.
Figure 7: Case study on segment ranking in the SY-Net514 network. Two highly ranked road segments (#470 and #336) identified by HetGL2R are highlighted in red to illustrate their structural positions in the network.
Table 9: Runtime Usage of HetGL2R and Baseline Methods on Three Networks.
@@
Figure 7 illustrates two representative road segments during ranking evaluation on the SY-Net514 network. In the rankings generated by HetGL2R, segment #470 has a relatively low traffic volume (215 vehicles/hour) but is ranked 15th, higher than segment #336, which carries a larger traffic volume (588 vehicles/hour, ranked 22nd). Further analysis shows that segment #470 is located near the destination of a certain OD pair, with upstream nodes carrying substantial traffic flow. A failure at segment #470 could block upstream traffic, causing congestion in the upstream region and potentially impacting the broader network performance. This example illustrates that HetGL2R accounts for upstream–downstream dependencies along traffic paths, beyond node attributes or local topological features.
@@ 5.4 Computational Efficiency and Scalability Analysis
Table 9 summarizes the per-epoch running time of HetGL2R on three networks with increasing scales. Compared with computationally lightweight methods (e.g., DeepWalk and TraNode2vec), specialized ranking models (e.g., LambdaLoss and RankFormer), and the ensemble model LightGBM, HetGL2R incurs higher training cost due to its more expressive architecture. However, when compared with other graph neural network–based approaches (e.g., PDFormer, DR-HGNN, and Edge-GNN), HetGL2R exhibits comparable computational efficiency. This efficiency can be attributed to its random-walk-based sampling strategy, which constrains computational complexity to scale linearly with walk length and sampling count. In contrast, models relying on global attention mechanisms or deep hierarchical message passing often incur quadratic or higher-order complexity with respect to the number of nodes. The primary scalability bottlenecks of HetGL2R stem from the Transformer encoder’s self-attention operations over sampled walk sequences and the storage and I/O overhead associated with full-graph walk sampling. As a result, while HetGL2R may not match the efficiency of lightweight models that do not exploit graph structure, it is expected to scale more favorably than most GNN-based methods on large graphs. Overall, by trading sampling for computation, HetGL2R achieves a practical balance between modeling expressiveness and computational feasibility.
@@ 5.5 Ablation Study
Ablation Study on HetGWalk. We evaluated the impact of joint random walks on the Trip Graph (TG) and Attribute-guided Graph (AG) by comparing the full model with a variant excluding AG (w/o AG). The w/o AG variant consistently shows substantial performance degradation across all networks and metrics, indicating that joint random walks are crucial for accurate and robust ranking. Using TG alone captures OD-induced functional dependencies along the graph structure, but does not explicitly model functional similarity, which can bias sampling toward high-centrality nodes and underexplore latent functional clusters. By enabling transitions between semantically similar nodes, AG enriches contextual diversity. Consequently, the joint TG–AG random walk integrates OD-induced functional dependence and semantic similarity, producing more informative sequences for robust segment representation.
Ablation on Transformer Encoder. We evaluated the role of the Transformer encoder by replacing it with GRU (Cho et al., 2014) and LSTM (Graves, 2012), denoted as w/o Transformer (GRU) and w/o Transformer (LSTM), respectively. The w/o Transformer variants consistently show inferior performance across all datasets and metrics, indicating that the Transformer encoder plays an important role in HetGL2R. RNN-based encoders rely on sequential propagation, which makes modeling long-range dependencies less effective in random-walk sequences. In contrast, the Transformer’s global self-attention mechanism enables direct interactions between arbitrary positions in a sequence, supporting more effective modeling of path-level dependencies and contributing to more robust ranking performance.
Ablation Study on AMIL Aggregation. We examined the impact of different context aggregation strategies by replacing AMIL with Mean, DeepSets (Zaheer et al., 2017), and SetTrans (Lee et al., 2019), denoted as w/o AMIL (Mean), w/o AMIL (DeepSets), and w/o AMIL (SetTrans), respectively. The variants without AMIL consistently show inferior performance across datasets and metrics, indicating the effectiveness of attention-based aggregation in HetGL2R. Mean aggregation treats all contexts equally and may be less effective at distinguishing critical contexts, while DeepSets aggregates instances independently and does not explicitly model inter-instance interactions. SetTrans, although expressive, may place excessive emphasis on fine-grained variations, making it more sensitive to noise. In contrast, AMIL employs a lightweight attention mechanism to selectively emphasize informative contexts while suppressing irrelevant ones, achieving a better balance between robustness and representation quality.
Ablation Study on KL-Divergence Loss. We examined the effect of different listwise ranking losses by replacing the KL-divergence loss with the cross-entropy loss used in ListNet (Cao et al., 2007) and the likelihood-based loss used in ListMLE (Xia et al., 2008a), denoted as w/o KL (ListNet) and w/o KL (ListMLE), respectively. The variants without KL divergence consistently show inferior performance across datasets and metrics, indicating the effectiveness of KL divergence for this task. This performance gap can be attributed to differences in modeling assumptions among listwise losses. ListMLE relies on stepwise ranking generation, which may accumulate errors when ranking supervision is uncertain, while ListNet focuses primarily on the Top-one probability and does not fully utilize ranking information from the remaining positions. In contrast, KL divergence performs a soft alignment between predicted and ground-truth importance distributions, emphasizing top-ranked segments while leveraging supervision from all positions, without relying on stepwise generation assumptions.
Figure 8: Parameter sensitivity analysis on three datasets. Each sub-plot shows the effect of a specific hyperparameter on model performance. The horizontal axis corresponds to the variation of the mixing coefficient $\alpha$, sampling sequence length lenlen, list length $K$, and depth-first exploration probability $\beta$, respectively, while the vertical axis represents the Kendall’s $\tau$ score.
@@
Overall, these ablation results suggest that each core component in the HetGL2R framework contributes meaningfully to the final performance. Their combination enables a balanced trade-off between modeling expressiveness and robustness for critical road segment identification.
@@ 5.6 Sensitivity analysis
We conducted a sensitivity analysis on key hyperparameters, including the walk bias factor $\alpha$, the sampling sequence length lenlen, the list length $K$, and the depth-first exploration probability $\beta$, with results summarized in Figure 8. When $\alpha$ is set to 0.6, the model achieves the highest Kendall’s $\tau$, suggesting a favorable trade-off between topological and attribute-related information. The optimal sampling length lenlen is 20, as shorter sequences provide limited contextual coverage, while longer sequences may introduce redundant or noisy information that degrades performance. Similarly, a list length of $K=5$ yields the best performance, as shorter lists offer insufficient ranking context, whereas longer lists increase task complexity and dilute effective ranking signals. Finally, setting $\beta$ to 0.8 results in the highest Kendall’s $\tau$, indicating that a moderate preference for depth-first exploration is beneficial. Smaller $\beta$ values underexploit multi-hop context, while $\beta=1.0$ shifts the walk toward a breadth-first–dominated exploration pattern, weakening deep relational signals and leading to noticeable performance degradation. Overall, these results demonstrate that HetGL2R is sensitive to key hyperparameters and provide empirical guidance for practical parameter selection.
5.7 Discussion
5.7.1 Analysis of Model Performance
The observed performance differences across baselines can be attributed to differences in their modeling assumptions and inductive biases, particularly in how structural cues, long-range dependencies, and path-based spatial influence are represented. Topology-dependent methods such as DeepWalk and IPageRank primarily rely on local connectivity patterns and centrality-related signals. These assumptions are often reasonable in structured urban networks, where network topology and OD flow are closely correlated. However, when this coupling weakens, their effectiveness degrades. This helps explain the pronounced performance drop observed in RD-Net3478, where network connectivity is generated independently of OD flow patterns, and highly connected or centrally located segments do not necessarily correspond to traffic corridors carrying substantial OD flow. Message-passing-based GNNs, including GAT and Dir-GNN, are further constrained by limited receptive fields and locality of aggregation. Even with multiple layers, information propagation is largely restricted to physical adjacency, making it challenging to capture long-range, path-dependent influence patterns induced by OD flows. More advanced heterogeneous models such as HGT and SeHGNN exhibit improved robustness on structured networks by incorporating type-aware aggregation and extended meta-paths. However, their reliance on message passing limits their ability to exploit sequential path semantics, such as node order and directional influence along OD paths, which are important for modeling traffic directionality and downstream effects. Notably, PDFormer achieves consistently strong performance across all three datasets. As a traffic-oriented model, it leverages semantic-mask-guided attention to capture long-range dependencies between geographically distant yet functionally related nodes and incorporates propagation-delay-aware mechanisms. While conceptually aligned with HetGL2R in emphasizing functional and long-range dependencies, PDFormer models such dependencies implicitly through masked attention over spatiotemporal views, which may be limited by error accumulation in similarity measurement (e.g., DTW-based semantic masking) over long distances due to signal attenuation and local noise interference. In contrast, HetGL2R explicitly represents OD flows, paths, and segment-level interactions, enabling a more fine-grained and explicit representation of traffic-induced influence propagation. Graph-based large language model approaches, such as ENGINE and GNNLLMEmb, exhibit limited effectiveness in this task. Although they provide rich semantic priors, they are designed for general graph structures and textual relational reasoning. Without explicit mechanisms to encode OD flows, path dependencies, or traffic-specific propagation dynamics, these models struggle to align their representations with the underlying mechanisms governing spatial influence in transportation networks. Overall, these observations suggest that the superior performance of HetGL2R is not solely due to the inclusion of OD information, but also to its sequence-based representation of spatial influence along traffic paths, which better aligns with the functional mechanisms underlying road-segment criticality.
5.7.2 Theoretical Implications
Building on the analysis above, we distill the broader theoretical insights. First, our findings provide empirical evidence that functional structure—induced by OD flows and paths—plays a more influential role than physical topology in modeling long-range spatial dependencies in road networks. This observation challenges the conventional reliance on purely topological measures and suggests that OD-aware modeling is important for understanding traffic-induced spatial influence. Second, by reframing road segment ranking as a learning-to-rank problem on heterogeneous graphs, our work introduces an alternative methodological lens for transportation network analysis. This formulation enables the explicit integration of OD flows and path structures—elements that are difficult to incorporate into standard graph neural network architectures—within a unified, task-driven learning framework. Third, our work offers an alternative perspective on the notion of ”importance” in networked systems. Unlike static centrality measures that treat importance as an inherent node property, our framework conceptualizes importance as context-dependent, shaped by demand patterns and usage paths.
5.7.3 Practical Implications
From a practical perspective, the resulting road segment rankings are intended to serve as a decision-support signal, providing an initial prioritization of potentially critical segments that can be combined with domain-specific operational objectives and policy constraints. Possible application scenarios include:
  • 1.
    Resilience assessment and monitoring. Transportation agencies can use the rankings to help identify segments whose failure may induce disproportionate network-wide disruption, supporting targeted monitoring and proactive reinforcement of high-risk corridors.
  • 2.
    Emergency response planning. By highlighting segments that are critical to high-demand OD pairs, the rankings can inform the design of contingency routing strategies that reduce reliance on vulnerable corridors during emergencies such as natural disasters or major incidents.
  • 3.
    Infrastructure investment prioritization. Under limited maintenance and upgrade budgets, the rankings can support the prioritization of road segments with higher potential systemic impact, complementing existing engineering and policy-based decision criteria.
  • 4.
    Transferability to other infrastructure networks. Although demonstrated on road networks, the core modeling principle—capturing functional dependencies induced by OD flows—is conceptually transferable to other infrastructure systems, such as power grids (electricity demand between substations) or communication networks (data flows between routers).
Together, these examples illustrate how the proposed framework can support practical decision-making by providing structured, system-level insights, while final actions remain guided by domain expertise and operational constraints.
@@ 6 Conclusion and Future Work
While the experimental evaluation is conducted on simulation-based networks, this choice reflects practical constraints in real-world traffic analysis, where large-scale controlled disruptions involve substantial cost, time, and institutional barriers. Simulation therefore provides a feasible environment for quantifying system-level criticality. Looking ahead, future work will focus on transferring the proposed framework to real-world settings by improving robustness to noisy or incomplete OD demand estimates, incorporating uncertainty-aware modeling, and validating the approach on large-scale operational networks as suitable data become available.